OpenStudy (anonymous):

Which point lies on the line described by the equation below? y + 4 = 4(x - 3)

OpenStudy (anonymous):

(2,-8)

OpenStudy (anonymous):

Check each of the given points. Substitute the x- and y- values into the equation. (x, y) Ex. Let's check if (2, -8) could work. (2, -8) (x, y) y + 4 = 4(x - 3) -8 + 4 = 4(2 - 3) Solve the equation. -8 + 4 = 4(2- 3) Distribute. Solve. -8 + 4 = 8 - 12 -4 = -4 ^---So this point could work. Check each of your options.